(x-2)(x+3)=(x-4)(x-5)
Answer by stanbon(48571)

(x-2)(x+3)=(x-4)(x-5)
x^2+x-6 = x^2-9x+20
x-6 = -9x+20
10x = 26
x = 2.6
--------------
Cheers,
Stan H.